OpenStudy (anonymous):

If the sum of three consecutive integers is 267. And your trying to find the largest integer, is 90 the correct answer? I want to make sure I'm right.

myininaya (myininaya):

let first integer be called n next integer is n+1 next next integer (n+1)+1=n+2 so we have n+(n+1)+(n+2)=267 3n+3=267 3(n+1)=267 n+1=267/3 n=267/3-1

myininaya (myininaya):

so n=89-1=88 so n+1=88+1=89 n+2=88+2=90

myininaya (myininaya):

yes 90 is right!

OpenStudy (anonymous):

Let the three numbers by \[x-1\]\[x\]\[x+1\] Adding the three we have, \[3x=267\] \[x=89\] So the three numbers are 88, 89, and 90 Hence you are right :) Cheers to that!

OpenStudy (anonymous):

thatnksss guysss:D